OpenStudy (anonymous):

How do I simplify 15z/8z-2

OpenStudy (anonymous):

Is the question\[\frac{ 15z }{ 8z } - 2\] or\[\frac{ 15z }{ 8z-2 } ?\]

jagr2713 (jagr2713):

@welcome to openstudy @anikkie

OpenStudy (anonymous):

it's the second one I'm new to this stuff so don't know yet how to actually write it

OpenStudy (jhannybean):

It's already simplied if it is \(\dfrac{15z}{8z-2}\)

jagr2713 (jagr2713):

yea

OpenStudy (anonymous):

oh ok thanks

OpenStudy (jhannybean):

Although you can pull out a 2, which is the LCM between 8 and 2, but that wont do anything.

OpenStudy (jhannybean):

\[\frac{15z}{2(4z-1)}\]

OpenStudy (anonymous):

Ok I see that

OpenStudy (jhannybean):

Mmhmm :)

OpenStudy (anonymous):

another problem I had was -20z-20 over 4z-1. so I simplified it to -5z-20 over z. is that correct?

OpenStudy (jhannybean):

\[\frac{-20z-20}{4z-1} = \frac{-20(z+1)}{4z-1}\]

OpenStudy (jhannybean):

You can only pull out the common terms between the numerator and denominator, therefore \(-5z-20\) is incorrect.

OpenStudy (anonymous):

omg how did you get the (z+1)

OpenStudy (jhannybean):

Pull out a common multiple of \(-20\) out from both values.

OpenStudy (anonymous):

I understand where the z came from but not the one

OpenStudy (jhannybean):

\(-20(z)=-20z~,~ -20(1)=-20 \therefore -20(z+1)=-20z-20\)

OpenStudy (anonymous):

got it im so bad at math and the thinking process for it. thanks

OpenStudy (jhannybean):

No problem. Any others?

OpenStudy (anonymous):

not yet

OpenStudy (jhannybean):

Alright. Close this question, and when you have another problem you need help solving,open up a new thread :D

OpenStudy (jhannybean):

Good luck.

OpenStudy (anonymous):

ok sure will